Shift Summary - Sept 13 1999 Day

Douglas A. Jensen (djensen@fnal.gov)
Mon, 13 Sep 1999 15:57:36 -0500 (CDT)

Shift Summary for DAY Shift, 09/13/99

On Shift: E. Munnier, D. Jensen

Hours of beam available at all intensities: 6.5h/8.5
Hours of beam available at nominal intensity: 8.5/8.8
Hours of E832 data (special runs not included): 6.5/8.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
It all started with the air conditioning down stairs allowing the
temperature to rise 2-4 deg. That killed off the HCC supplies.
Mucking with the air conditioner generates various temperature
excursions. Had turned up temperature threshold about 3 deg.
Then, in sympathy, the Regenerator HV started tripping. Ed and
Elizabeth go down and unplug and replug interlock cables to fix
the problem. Shift ends with us still in access mode, waiting
to see if regenerator HV stays on, and air conditioning guys
still mucking - recharging west side compressors. There was also
a lot of ice buildup on some of the coils - indicitative of needing
a recharge.

========================================================================

Runs ended this shift:

14512 BEAM_832 22.5 M events

Tapes brought to FCC this shift:


========================================================================

Accesses during shift:
Sydney and Doug - turn on ET power suplly
Ed and Elizabeth - work on Regenerator HV tripping problem.

========================================================================

Overview of shift: 832 data taking at 4.0 E12
Reminder ET3 is at 28.4 instead of the expected 25 GeV. Sydney says OK
As per Ed request, Sydney called about split disks filling. Not critical
yet. Then he solved the problem.
In the DC summary timing plots, all the plots have less background.
Air conditioning problems for the last part of the shift.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14512 294 8:20 KFREND
Leonard Nelson stopped by. He is chasing down
maint., etc of air systems. Discuss with FESS etc.
9:05 CsI channel 2414, as in prev shift, at 4.46% off
in laser stability.
329 9:08 CsI Panic message - goes away. Problem with a
channel at 4.4% also disappears.
345 Channel 1668 at about 4.7% on laser, but disappears
back into the peak after a few spills.
390 10:30 KFRIEND
423 11:15 KFRIEND
noon Mounted QKNL35 as per request.
13:25 Two new He tubes on line, DN says that he
will add alcohol tomorrow.
524 13:30 HCC Inefficient messages from Elliot
Turns out - air conditioning problems leading to
ET power supplies triping. That grinds us to
a halt for a while.
14:05 End run, exchange tapes etc.
Started and aborted a couple of runs because of
regenerator HV tripping off.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

If humidity alarms go off (humidity above 10% in the blockhouse)
call Leonard Nelson. His home number is 630-231-1416; pager
888-750-0702.

Note:
Ed reports from 9 am meeting that the beam will go off Wed AM at
5:30 AM. as scheduled.

STILL! No working 'Xerox' Machine. My it has been a long time!